OpenStudy (anonymous):

find the inverse of f(x) = x+3/2

OpenStudy (nottim):

Switch x and y.

OpenStudy (nottim):

x=y+3/2

OpenStudy (nottim):

Do you know what to do after?

OpenStudy (anonymous):

no

OpenStudy (nottim):

TELL ME WHERE THE DRUGS ARE

OpenStudy (nottim):

Ok. Er. You don't got a textbook?

OpenStudy (nottim):

Anyways.

OpenStudy (nottim):

How do you isolate y now?

OpenStudy (nottim):

FEAR ME

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

x-y=3/2?

OpenStudy (nottim):

What?

OpenStudy (nottim):

Isolate y, not give it x as a buddy.

OpenStudy (anonymous):

so u subtract x-3/2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then whats next

OpenStudy (nottim):

That's it. Congrats, YOU SOLVED IT BY YOURSELF!

OpenStudy (anonymous):

can u help me with more pls

OpenStudy (anonymous):

how do you solve the inequality x^2<2

OpenStudy (anonymous):

can you help me with my question?

OpenStudy (nottim):

ER, post a new question? I'll answer there.

OpenStudy (nottim):

But its the same proces yo udid earlier. Just algebra.